# Split Unlearningthanks: Accepted by ACM CCS’25

### v6

@@ Appendix D Evaluating Sample-Level Unlearning Tasks
Consistent with the setup of the previous experiments in Section 6.3, we utilize the simplified-VGG [46] and ResNet18 [47] as the model architectures for the split (un)learning experiments, considering four different data distribution scenarios across three datasets.
@@
Additionally, we examine the impact on membership inference attacks targeting the unlearned samples $\mathbb{X}_{u}$. The success probability of performing membership inference against the requesting client $\mathbb{C}_{u}$ is measured using the hit rate of $\mathbb{X}_{u}$. As shown in Table 11, our frameworks significantly reduce the hit rate of $\mathbb{X}_{u}$, thereby exhibiting stronger resistance to membership inference attacks [34, 56] for $\mathbb{C}_{u}$. Compared to Vanilla SL, our frameworks provide enhanced protection against such attacks, particularly in SplitWiper+, where the hit rate of $\mathbb{X}_{u}$ is further minimized, demonstrating superior resilience to member inference vulnerabilities.
@@
Clients with same task Partially overlapping labels AG-News L. $\mathbb{C}_{u}$ 34.77/51.92 33.98/52.08 1.34/1.74 3.71/3.44 5.80/5.52 2.95/3.02 $\mathbb{Y}_{u}$ 81.94 88.47 89.21 90.31 90.89 90.37 90.68 91.16 92.32
$\mathbb{C}_{o}$ 30.55/49.26 30.30/49.80 1.03/1.65 3.25/3.26 5.12/4.86 2.81/3.05 $\mathbb{Y}_{o}$ 83.14 85.68 85.16 86.24 86.67 85.28 85.99 85.08 86.41
U. $\mathbb{C}_{u}$ 16.45/40.47 16.16/40.56 0.53/1.35 1.78/2.57 3.01/4.39 1.76/2.26 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 30.25/49.03 32.57/49.44 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 82.21 85.19 85.82 86.46 86.71 86.07 86.47 87.19 87.32
@@
Clients with same task Partially overlapping labels AG-News L.   82.95/82.19 85.68/89.08 85.54/89.39 85.77/90.12 86.89/91.10 85.04/90.89 85.69/90.50 85.47/91.11 86.39/91.74 BERT
@@
Clients with same task Partially overlapping labels AG-News  Learning   97.32 99.97 99.96 99.79 99.86 99.90 99.99 99.97 99.97 BERT
@@
Clients with same task Homo labels Hetero volumes CIFAR-10 L.   77.04/90.30 78.08/88.60 79.86/89.60 80.65/89.10 81.08/90.05 80.4/88.20 80.82/90.10 80.73/89.30 81.06/90.80 Resnet
@@
0.9   66.88/17.10 67.13/42.00 71.62/43.00 70.77/40.30 71.08/45.60 67.67/37.50 69.53/41.40 70.80/43.60 72.04/39.80
Partially overlapping labels CIFAR-10 L.   73.30/85.20 78.59/86.80 80.91/88.10 80.34/88.90 80.39/90.70 80.57/87.90 80.11/88.60 80.92/87.40 81.31/89.30
@@
0.9   63.86/18.30 63.43/48.10 69.92/50.20 68.93/47.50 69.23/47.30 68.29/42.70 68.74/40.95 70.03/42.70 71.02/51.05
Hetero labels CIFAR-10 L.   85.66/88.60 85.65/86.70 87.03/92.10 87.89/91.30 88.36/92.90 86.66/91.40 87.77/92.30 88.13/91.80 88.76/93.30
@@
0.9   69.85/18.60 70.81/47.40 74.08/53.00 71.43/44.50 73.08/45.10 71.67/50.10 72.22/46.95 72.81/48.90 73.23/47.30
different tasks CIFAR-10 + MNIST L.   89.57/91.50 88.62/90.60 94.07/93.40 93.19/91.90 93.92/92.80 92.66/91.40 92.35/92.30 93.36/92.60 93.40/92.40
@@
0.9   90.49/28.40 90.58/43.20 94.35/46.70 93.74/38.00 93.56/39.90 93.06/42.30 92.98/42.30 93.60/43.00 93.52/43.20
Clients with same task Homo labels Hetero volumes CIFAR-10 L.   93.95/94.70 92.35/92.50 96.83/99.50 96.56/98.30 97.49/98.20 96.64/97.60 95.75/97.40 96.64/98.50 97.22/99.30 VGG
@@
0.9   82.34/37.10 81.97/60.20 86.27/57.00 83.76/47.10 84.69/48.90 82.35/46.10 84.21/44.70 86.03/51.30 87.04/50.50
Partially overlapping labels CIFAR-10 L.   94.25/95.80 92.23/93.70 97.57/99.20 96.73/98.70 97.00/98.80 95.53/99.20 96.85/99.60 96.91/99.20 97.75/99.30
@@
0.9   82.53/30.00 82.06/60.30 85.59/62.90 83.69/49.60 84.51/57.10 83.95/40.30 84.03/43.00 83.97/57.70 84.95/59.10
Hetero labels CIFAR-10 L.   92.78/93.70 93.78/93.20 98.98/99.40 97.44/99.70 98.11/99.80 95.96/99.80 96.73/99.50 97.81/99.90 98.47/99.90
@@
0.9   82.19/23.00 80.50/59.10 88.64/58.60 86.06/48.00 87.85/53.50 84.31/44.40 86.43/47.00 87.40/59.00 87.72/56.40
different tasks CIFAR-10 + MNIST L. 94.95/95.40 95.31/95.20 98.08/99.90 96.06/97.70 97.10/98.80 95.13/97.10 96.90/97.00 96.52/97.80 97.11/97.90
@@ Appendix E Experiments on Text Modality: AG-News with BERT
To further demonstrate the generality of our framework, SplitWiper and SplitWiper+, we extend our evaluation beyond vision tasks. Unlike methods tailored to specific datasets or architectures, our frameworks are designed to support a wide range of data modalities and can be seamlessly integrated with state-of-the-art models. While Section 6 presents strong performance on standard image benchmarks such as CIFAR-10, CIFAR-100, and MNIST, we supplement our analysis with experiments on a non-vision dataset: AG-News [57], using the mini-BERT [58] architecture.
@@ Appendix F Experiments on Label Privacy under BIRCH-based Inference Attacks
To further validate the robustness of SplitWiper+ in protecting label privacy and the privacy of intermediate representations, we extend our evaluation beyond K-Means by introducing BIRCH [59]. BIRCH is a hierarchical clustering method and provides a complementary perspective with different clustering assumptions. This allows us to better assess whether SplitWiper+ can defend against a broader class of unsupervised inference attacks.
@@
Clients with same task Homogeneous labels, heterogeneous volumes CIFAR-10  Learning   86.94 97.48 90.58 89.42 90.62 88.68 89.86 90.54 90.88 Resnet
@@
Partially overlapping labels CIFAR-10 Learning   86.14 99.34 89.72 90.92 90.88 88.46 89.16 90.68 90.58
@@
Heterogenous labels CIFAR-10 Learning   86.24 98.70 97.52 91.54 90.32 91.20 90.28 93.54 92.44
@@
different tasks CIFAR-10 + MNIST Learning   96.76 98.96 98.31 92.20 90.10 90.72 88.12 90.62 88.62
@@
Clients with same task Homogeneous labels, heterogeneous volumes CIFAR-10 Learning   88.10 99.20 99.14 98.60 96.92 94.50 95.08 95.54 96.74 VGG
@@
Partially overlapping labels CIFAR-10 Learning   92.04 99.80 99.22 95.56 98.56 94.08 98.60 97.86 98.56
@@
Heterogeneous labels CIFAR-10 Learning   90.92 99.88 99.60 94.02 97.58 99.40 96.12 98.66 98.45
@@
different tasks CIFAR-10 + MNIST Learning   89.54 98.84 98.38 97.86 96.78 93.38 96.27 96.26 96.36
@@ Appendix G Evaluating Convergence
Establishing theoretical proof of convergence is challenging [13] for SplitWiper and SplitWiper+. Therefore, we rely on empirical experiments to assess convergence by tracking the number of nodes that change in each iteration. The experimental data indicates that the proportion of moved nodes steadily approaches zero within 30 epochs when using the CIFAR-10 dataset or a combination of CIFAR-10 and MNIST; for CIFAR-100, this occurs within 50 epochs.

### v8

@@ Appendix D Evaluating Sample-Level Unlearning Tasks
Consistent with the setup of the previous experiments in Section 6.3, we utilize the simplified-VGG [bao2024efficient] and ResNet18 [he2016deep] as the model architectures for the split (un)learning experiments, considering four different data distribution scenarios across three datasets.
@@
Additionally, we examine the impact on membership inference attacks targeting the unlearned samples $\mathbb{X}_{u}$. The success probability of performing membership inference against the requesting client $\mathbb{C}_{u}$ is measured using the hit rate of $\mathbb{X}_{u}$. As shown in Table 11, our frameworks significantly reduce the hit rate of $\mathbb{X}_{u}$, thereby exhibiting stronger resistance to membership inference attacks [10.1145/3460120.3484756, lu2022label] for $\mathbb{C}_{u}$. Compared to Vanilla SL, our frameworks provide enhanced protection against such attacks, particularly in SplitWiper+, where the hit rate of $\mathbb{X}_{u}$ is further minimized, demonstrating superior resilience to member inference vulnerabilities.
@@
Clients with same task
Partially overlapping labels
AG-News
L.
$\mathbb{C}_{u}$
34.77/51.92 33.98/52.08 1.34/1.74 3.71/3.44 5.80/5.52 2.95/3.02 $\mathbb{Y}_{u}$ 81.94 88.47 89.21 90.31 90.89 90.37 90.68 91.16 92.32
$\mathbb{C}_{o}$
30.55/49.26 30.30/49.80 1.03/1.65 3.25/3.26 5.12/4.86 2.81/3.05 $\mathbb{Y}_{o}$ 83.14 85.68 85.16 86.24 86.67 85.28 85.99 85.08 86.41
U.
$\mathbb{C}_{u}$
16.45/40.47 16.16/40.56 0.53/1.35 1.78/2.57 3.01/4.39 1.76/2.26 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$
30.25/49.03 32.57/49.44 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 82.21 85.19 85.82 86.46 86.71 86.07 86.47 87.19 87.32
@@
Clients with same task
Partially overlapping labels
AG-News
L.   82.95/82.19 85.68/89.08 85.54/89.39 85.77/90.12 86.89/91.10 85.04/90.89 85.69/90.50 85.47/91.11 86.39/91.74 BERT
@@
Clients with same task
Partially overlapping labels
AG-News
 Learning   97.32 99.97 99.96 99.79 99.86 99.90 99.99 99.97 99.97 BERT
@@
Clients with same task
Homo labels Hetero volumes
CIFAR-10
L.   77.04/90.30 78.08/88.60 79.86/89.60 80.65/89.10 81.08/90.05 80.4/88.20 80.82/90.10 80.73/89.30 81.06/90.80 Resnet
@@
0.9
  66.88/17.10 67.13/42.00 71.62/43.00 70.77/40.30 71.08/45.60 67.67/37.50 69.53/41.40 70.80/43.60 72.04/39.80
Partially overlapping labels
CIFAR-10
L.   73.30/85.20 78.59/86.80 80.91/88.10 80.34/88.90 80.39/90.70 80.57/87.90 80.11/88.60 80.92/87.40 81.31/89.30
@@
0.9
  63.86/18.30 63.43/48.10 69.92/50.20 68.93/47.50 69.23/47.30 68.29/42.70 68.74/40.95 70.03/42.70 71.02/51.05
Hetero labels
CIFAR-10
L.   85.66/88.60 85.65/86.70 87.03/92.10 87.89/91.30 88.36/92.90 86.66/91.40 87.77/92.30 88.13/91.80 88.76/93.30
@@
0.9
  69.85/18.60 70.81/47.40 74.08/53.00 71.43/44.50 73.08/45.10 71.67/50.10 72.22/46.95 72.81/48.90 73.23/47.30
different tasks
CIFAR-10 + MNIST
L.   89.57/91.50 88.62/90.60 94.07/93.40 93.19/91.90 93.92/92.80 92.66/91.40 92.35/92.30 93.36/92.60 93.40/92.40
@@
0.9
  90.49/28.40 90.58/43.20 94.35/46.70 93.74/38.00 93.56/39.90 93.06/42.30 92.98/42.30 93.60/43.00 93.52/43.20
Clients with same task
Homo labels Hetero volumes
CIFAR-10
L.   93.95/94.70 92.35/92.50 96.83/99.50 96.56/98.30 97.49/98.20 96.64/97.60 95.75/97.40 96.64/98.50 97.22/99.30 VGG
@@
0.9
  82.34/37.10 81.97/60.20 86.27/57.00 83.76/47.10 84.69/48.90 82.35/46.10 84.21/44.70 86.03/51.30 87.04/50.50
Partially overlapping labels
CIFAR-10
L.   94.25/95.80 92.23/93.70 97.57/99.20 96.73/98.70 97.00/98.80 95.53/99.20 96.85/99.60 96.91/99.20 97.75/99.30
@@
0.9
  82.53/30.00 82.06/60.30 85.59/62.90 83.69/49.60 84.51/57.10 83.95/40.30 84.03/43.00 83.97/57.70 84.95/59.10
Hetero labels
CIFAR-10
L.   92.78/93.70 93.78/93.20 98.98/99.40 97.44/99.70 98.11/99.80 95.96/99.80 96.73/99.50 97.81/99.90 98.47/99.90
@@
0.9
  82.19/23.00 80.50/59.10 88.64/58.60 86.06/48.00 87.85/53.50 84.31/44.40 86.43/47.00 87.40/59.00 87.72/56.40
different tasks
CIFAR-10 + MNIST
L. 94.95/95.40 95.31/95.20 98.08/99.90 96.06/97.70 97.10/98.80 95.13/97.10 96.90/97.00 96.52/97.80 97.11/97.90
@@ Appendix E Experiments on Text Modality: AG-News with BERT
To further demonstrate the generality of our framework, SplitWiper and SplitWiper+, we extend our evaluation beyond vision tasks. Unlike methods tailored to specific datasets or architectures, our frameworks are designed to support a wide range of data modalities and can be seamlessly integrated with state-of-the-art models. While Section 6 presents strong performance on standard image benchmarks such as CIFAR-10, CIFAR-100, and MNIST, we supplement our analysis with experiments on a non-vision dataset: AG-News [zhang2015character], using the mini-BERT [dong2023rai2] architecture.
@@ Appendix F Experiments on Label Privacy under BIRCH-based Inference Attacks
To further validate the robustness of SplitWiper+ in protecting label privacy and the privacy of intermediate representations, we extend our evaluation beyond K-Means by introducing BIRCH [zhang1996birch]. BIRCH is a hierarchical clustering method and provides a complementary perspective with different clustering assumptions. This allows us to better assess whether SplitWiper+ can defend against a broader class of unsupervised inference attacks.
@@
Clients with same task
Homogeneous labels, heterogeneous volumes
CIFAR-10
 Learning   86.94 97.48 90.58 89.42 90.62 88.68 89.86 90.54 90.88 Resnet
@@
Partially overlapping labels
CIFAR-10
Learning   86.14 99.34 89.72 90.92 90.88 88.46 89.16 90.68 90.58
@@
Heterogenous labels
CIFAR-10
Learning   86.24 98.70 97.52 91.54 90.32 91.20 90.28 93.54 92.44
@@
different tasks
CIFAR-10 + MNIST
Learning   96.76 98.96 98.31 92.20 90.10 90.72 88.12 90.62 88.62
@@
Clients with same task
Homogeneous labels, heterogeneous volumes
CIFAR-10
Learning   88.10 99.20 99.14 98.60 96.92 94.50 95.08 95.54 96.74 VGG
@@
Partially overlapping labels
CIFAR-10
Learning   92.04 99.80 99.22 95.56 98.56 94.08 98.60 97.86 98.56
@@
Heterogeneous labels
CIFAR-10
Learning   90.92 99.88 99.60 94.02 97.58 99.40 96.12 98.66 98.45
@@
different tasks
CIFAR-10 + MNIST
Learning   89.54 98.84 98.38 97.86 96.78 93.38 96.27 96.26 96.36
@@ Appendix G Evaluating Convergence
Establishing theoretical proof of convergence is challenging [chen2022graph] for SplitWiper and SplitWiper+. Therefore, we rely on empirical experiments to assess convergence by tracking the number of nodes that change in each iteration. The experimental data indicates that the proportion of moved nodes steadily approaches zero within 30 epochs when using the CIFAR-10 dataset or a combination of CIFAR-10 and MNIST; for CIFAR-100, this occurs within 50 epochs.